STATE-RESOLVED PREDISSOCIATION DYNAMICS OF THE FORMYLOXYL RADICAL BY DISSOCIA-TIVE PHOTODETACHMENT OF HCO_{2}^{−} / DCO_{2}^{−}

\underline{AMELIA W. RAY}, BEN B. SHEN, BERWYCK L. J. POAD and ROBERT E. CONTINETTI^{a}, Department of Chemistry and Biochemistry, University of California, San Diego, La Jolla, CA 92093.

Near threshold dissociative photodetachment dynamics of HCO_{2}^{−} and DCO_{2}^{−} were studied using photoelectronphotofragment coincidence spectroscopy in a cryogenically cooled ion beam trap. The combination of colder precursor ions and lower kinetic energy electrons has enabled us to record higher resolution state-resolved dissociation dynamics for the HCO_{2}^{−}/DCO_{2}^{−} + hν → H/D + CO_{2} + e^{−} reaction, complementing and extending our earlier study of this system.^{b} For hν = 4.27 eV, photodetachment accesses the three lowest-lying electronic states (^{2}B_{2}, ^{2}A_{1}, and ^{2}A_{2}) of the formyloxyl radical, with the resulting vibrational features dominated by the OCO bending mode. Excitation of this OCO bend in the transient formyloxyl neutral couples to the bending excitation observed in the state-resolved translational energy distribution for the CO_{2} product. Each vibrational state of the nascent neutral is observed to predissociate to several CO_{2} product states leading to pronounced vibrational predissociation sequence bands in the photoelectron-photofragment coincidence spectra for both HCO_{2}^{−} and DCO_{2}^{−}.

aThis work is supported by the US Department of Energy under Grant Number DE-FG03-98ER14879

bT.G. Clements and R.E. Continetti, J. Chem. Phys., 115, 5345 (2001)